Mathematics 242 – Principles of Analysis Definitions Quiz 5 – October 21, 2005

Your Name:

1. What is the definition of convergence for an infinite series ∑_{k=1}^{∞} a_{k}?

2. Give an example of a conditionally convergent series, and explain how you can tell that it is.